OpenStudy (anonymous):

PLEASE HELP!! A hair product company sells three types of hair products for $30, $20, and $10 per unit. In 1 year, the total revenue for the three products was $800,000, which corresponded to the sale of 40,000 units. The company sold half as many units of the $30 product as of the $20 product. How many units of each product were sold?

OpenStudy (anonymous):

Not really sure where to go from there? @robtobey

OpenStudy (campbell_st):

well you know that if you let the $20 product = 2a then the $30 product is a let the $10 product = b so set up the equations revenue 30a + 20(2a) + 10b = 800000 and units a + 2a + b = 40000 so the simultaneous equations are 70a + 10b = 800000 (1) 3a + b = 40000 (2) multiply equation (2) by 10 and subtract it from (1) 70a + 10b = 800000 - 30a + 10b = 400000 ---------------------- 40a = 400000 then a = 10000 so substitute this into equation 2 to find b 30000 + b = 40000 so b = 10000 then the company solve 10000 units of $30, 20000 units of $20 and 10000 units of $10 hope it helps